OpenStudy (anonymous):

Simplify: (3x + 12) (3x - 12).

OpenStudy (anonymous):

(a + b)(a - b) = a^2 - b^2 Here, a = 3x and b = 12 If you can compute (3x)^2 and (12)^2 , you will be able to get the rest of this answer.

OpenStudy (anonymous):

so 3x^2 - 12^2 ?

OpenStudy (anonymous):

One more hint: (cd)^2 = (c^2)(d^2) So, you have to square the "3" and the "x".

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

What is 3^2?

OpenStudy (anonymous):

3x squared

OpenStudy (anonymous):

3 squared i mean

OpenStudy (anonymous):

9

OpenStudy (anonymous):

3*3 is 9 i mean

OpenStudy (anonymous):

so 9x^2 then what

OpenStudy (anonymous):

3^2 is "3 squared". That means: 3 times 3. That is 9, so: (3x)^2 = 9x^2 not 3x^2 Also 12^2 = 144 So, 9x^2 - 144

OpenStudy (anonymous):

So, (3x + 12)(3x - 12) = 9x^2 - 144

OpenStudy (anonymous):

This is because:: (a + b)(a - b) = a^2 + ab - ab - b^2 = a^2 - b^2